k^-1(x) when k(x) = x^2+5
Answer by stanbon(75887)

k^-1(x) when k(x) = x^2+5
----
y = x^2+5
---
1st: Interchange x and y to get:
x = y^2+5
---
2nd: Solve for "y":
y^2 = x-5
y = sqrt(x-5)
====================
Cheers,
Stan H.